The proportions of blood phenotypes A, B, AB, and O in the population of all Caucasians in the united states are reported as .40, .11, .04, and .45 respectively. If a single Caucasian is chosen randomly from the population, what is the probability that he or she will have either type A or type AB blood?
Answer #1

Solution :

\mathrm{Given\::}

\mathrm{P\left(A\right)=0.40,\:\:P\left(B\right)=0.11,\:P\left(AB\right)=0.04\:\:and\:\:P\left(O\right)=0.45}

\mathrm{Now,}

\mathrm{Probability\:that\:he\:or\:she\:will\:have\:either\:type\:A\:or\:type\:AB\:blood\:=\:P\left(A\right)+P\left(AB\right)}

\mathrm{=0.40+0.04}

\mathrm{=0.44}